# A continuum mechanics approach for the deformation of non-Euclidean origami generated by piecewise constant nematic director fields

### v1

@@ 3 Three-fold intersection
In this section, we study the kinematics of three-fold intersections with patterned piecewise constant director fields satisfying the metric compatibility condition at each interface. Our approach is shown in Fig. 3(a). The reference pattern is metric compatible at each interface. To compute the actuated configuration, the first step is to cut the reference pattern along the interface ${\bf t}_{1}$ and then deform it by the spontaneous deformation gradient ${\bf U}_{{\bf n}_{i}}$ at each domain. This will form an intermediate configuration in 2D, with a deficit or a surplus. The second step is to deform the intermediate configuration isometrically and stitch the boundaries together to form a non-Euclidean origami with Gaussian curvature at the tip. We use $\alpha_{i}$ and $\tilde{\alpha}_{i}$ to denote the sector angles before and after actuation, respectively.
@@ 3.1 Three-fold origami with angular deficit and surplus
for both the deficit case and the surplus case. Satisfying the compatibility condition yields a folding deformation ${\bf y}(\tilde{{\bf x}})$ that maps the intermediate state to the folded state as
@@
$\gamma_{2}$ $=$ $\sigma\arccos\left(\frac{-\cos\tilde{\alpha}_{3}+\cos\tilde{\alpha}_{1}\cos\tilde{\alpha}_{2}}{\sin\tilde{\alpha}_{1}\sin\tilde{\alpha}_{2}}\right),$
$\gamma_{3}$ $=$ $\sigma\operatorname{sign}\left(\frac{\sin\tilde{\alpha}_{1}}{\sin\tilde{\alpha}_{3}}\right)\arccos\left(\frac{-\cos\tilde{\alpha}_{1}+\cos\tilde{\alpha}_{2}\cos\tilde{\alpha}_{3}}{\sin\tilde{\alpha}_{2}\sin\tilde{\alpha}_{3}}\right),$ (9)
for $\tilde{\alpha}_{i}$ in the admissible shown in Fig. 4(a), where $\sigma\in{\pm 1}$ denotes the two branches.
@@
${\bf R}_{3}(\gamma_{3})\tilde{{\bf t}}_{4}\cdot\tilde{{\bf t}}_{2}=\tilde{{\bf t}}_{1}\cdot\tilde{{\bf t}}_{2}\Leftrightarrow\cos\gamma_{3}=\frac{-\cos\tilde{\alpha}_{1}+\cos\tilde{\alpha}_{2}\cos\tilde{\alpha}_{3}}{\sin\tilde{\alpha}_{2}\sin\tilde{\alpha}_{3}}.$ (10)
@@
$\gamma_{2}$ $=$ $\sigma\arccos\left(\frac{-\cos\tilde{\alpha}_{3}+\cos\tilde{\alpha}_{1}\cos\tilde{\alpha}_{2}}{\sin\tilde{\alpha}_{1}\sin\tilde{\alpha}_{2}}\right),$
$\gamma_{3}$ $=$ $\sigma\operatorname{sign}\left(\frac{\sin\tilde{\alpha}_{1}}{\sin\tilde{\alpha}_{3}}\right)\arccos\left(\frac{-\cos\tilde{\alpha}_{1}+\cos\tilde{\alpha}_{2}\cos\tilde{\alpha}_{3}}{\sin\tilde{\alpha}_{2}\sin\tilde{\alpha}_{3}}\right),$ (13)
where $\sigma\in\{\pm 1\}$ denotes the two branches. Since $\cos\gamma_{2},\cos\gamma_{3}\in[-1,1]$, we have the admissible domain (Fig. 4(a)) for $\tilde{\alpha}_{1,2,3}$ such that the absolute values of Eqs. (10) and (11) are less than or equal to 1. ∎
Figure 4: (a) The admissible domain of $\tilde{\alpha}_{1}$, $\tilde{\alpha}_{2}$ and $\tilde{\alpha}_{3}$. (b) Notation for a three-fold director field.
@@ 3.2 Connection to director fields
where $\sigma_{i},\tilde{\sigma}_{i}\in\{\pm 1\}$. Equation (14) implies that there are four solutions of ${\bf t}_{2}$ for the prescribed ${\bf n}_{1}$ and ${\bf n}_{2}$. However, these tangents cannot be arbitrarily chosen from the four solutions when patterning a three-fold pattern--they need to obey the following topological condition
@@
An interesting observation from Theorem 3.2 is that, fixing $\sigma_{i}$ and $\tilde{\sigma}_{i}$, the folding angle $\gamma_{i}$ depends only on the directors on its adjacent panels, i.e., ${\bf n}_{i}$ and ${\bf n}_{i-1}$ with the directional angles $\theta_{i}$ and $\theta_{i-1}$. This fact provides more freedom to design connected three-fold intersections. For example, in Fig. 5, the pattern consists of two three-fold intersections. We may keep ${\bf n}_{2}$ and ${\bf n}_{3}$ unchanged and vary the direction of ${\bf n}_{4}$. As a result of Theorem 3.2, the folding angle along the shared crease in the actuated configuration is independent of ${\bf n}_{4}$ and will remain unchanged. Thus, the actuated configuration containing the two three-fold nodes remains compatible. We will use the basic pattern in Fig. 5 to design more complex patterns in Section 5.
@@ 4 Four-fold intersection
In this section, we study the kinematics of a general non-Euclidean four-fold origami generated by piecewise constant director fields. Again, we deform the intermediate state isometrically to achieve a continuous folded state. The method we use to compute the folding angle functions is partially adapted from our previous work zou2024kinematics . Here, we use a new symmetry approach to make the folding angle functions rigorous and more convenient to use for computation. We also provide a more physically relevant understanding of the “branches” in Fig. 7.
@@ 4.1 Reference domain
Figure 6: (a) Two equivalent intermediate configurations. (b) Two folding branches denoted by $\sigma=+1$ and $\sigma=-1$.
[…]
@@ 4.2 Compatibility condition and folding angle functions
The folding angles are restricted in the interval $-\pi\leq\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4}\leq\pi$. In the rigidly and flat-foldable case, the four folding angles reach ±$\pm\pi$ simultaneously, with a reference free folding angle ranging from $-\pi$ to $\pi$. In the non-Euclidean case, some folding angles will reach ±$\pm\pi$ first, making the domain of the reference folding angle not a full interval of $[-\pi,\pi]$.
@@ 4.2.1 Calculations of $\gamma_{2}$, $\gamma_{3}$ and $\gamma_{4}$
$\gamma_{4}=\bar{\gamma}_{4}^{\sigma}(\theta):=\sigma\arccos F_{4}(\theta;\tilde{\boldsymbol{\alpha}}).$ (33)
Here $\sigma\in\{\pm\}$ denotes the two branches of solution as shown in Fig. 6(b). Since the function $\bar{\gamma}_{4}^{\sigma}(\theta)$ is even and monotonic in $[-\pi,0]$ and $[0,\pi]$, we may characterize the admissible domain of $\theta$ by ${\cal I}^{-}\cup{\cal I}^{+}:=[-\theta_{\max},-\theta_{\min}]\cup[\theta_{\min},\theta_{\max}]$ in the following cases:
@@
    Otherwise, if $F_{4}(0;\tilde{\boldsymbol{\alpha}})$ and $F_{4}(\pi;\tilde{\boldsymbol{\alpha}})$ are beyond the above range, the admissible set of $\theta$ is empty.
@@
${\bf R}_{3}(\gamma_{3})(({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{4}(\bar{\gamma}_{4}^{\sigma}(\gamma_{2}))\tilde{{\bf t}}_{5})=({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}.$ (37)
@@
which is trivially true. Thus, to solve the projected equation (37), we consider the non-degenerate case under $|({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}|\neq 0$. For (37) to hold, $\gamma_{3}=\bar{\gamma}_{3}^{\sigma}(\gamma_{2})$ must satisfy
$\cos(\bar{\gamma}_{3}^{\sigma}(\gamma_{2}))$ $=$ $\frac{({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{4}(\bar{\gamma}_{4}^{\sigma}(\gamma_{2}))\tilde{{\bf t}}_{5}\cdot({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}}{|({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{4}(\bar{\gamma}_{4}^{\sigma}(\gamma_{2}))\tilde{{\bf t}}_{5}||({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}|}$ (39)
::=:= $F_{3c}^{\sigma}(\gamma_{2};\tilde{\boldsymbol{\alpha}}),$
$\sin(\bar{\gamma}_{3}^{\sigma}(\gamma_{2}))$ $=$ $\text{sign}\left[\tilde{{\bf t}}_{3}\cdot({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{4}(\bar{\gamma}_{4}^{\sigma}(\gamma_{2}))\tilde{{\bf t}}_{5}\times({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}\right]\sqrt{1-\cos^{2}(\bar{\gamma}_{3}^{\sigma}(\gamma_{2}))}$ (40)
::=:= $F_{3s}^{\sigma}(\gamma_{2};\tilde{\boldsymbol{\alpha}}).$
@@
$\gamma_{3}=\bar{\gamma}_{3}^{\sigma}(\gamma_{2})=\text{sign}[F_{3s}^{\sigma}(\gamma_{2};\tilde{\boldsymbol{\alpha}})]\arccos[F_{3c}^{\sigma}(\gamma_{2};\tilde{\boldsymbol{\alpha}})].$ (41)
@@ 4.2.2 Calculation of $\gamma_{1}$
To solve $\gamma_{1}$, we use the second intermediate domain in Fig. 6(a) such that there exists $\gamma_{1}=\bar{\gamma}_{1}^{\sigma}(\theta)$ making the folded configuration continuous across $\tilde{{\bf t}}_{4}$ and $\tilde{{\bf t}}_{55}$. A convenient way to obtain $\gamma_{1}$ is to use the cyclic symmetry of $\tilde{\alpha}_{1},\tilde{\alpha}_{2},\tilde{\alpha}_{3}$ and $\tilde{\alpha}_{4}$. Notice that $\gamma_{4}=\bar{\gamma}_{4}^{\sigma}(\gamma_{2})$ shows the ability to solve the folding angle ($\gamma_{4}$) in terms of its opposite folding angle ($\gamma_{2}$). Thus we may solve $\gamma_{1}$ in terms of its opposite folding angle $\gamma_{3}$, by simply cycling $\tilde{\alpha}_{1},\tilde{\alpha}_{2},\tilde{\alpha}_{3},\tilde{\alpha}_{4}$ and then using the the same function form $F_{4}$ in (32) as follows:
$\gamma_{1}=\hat{\sigma}\arccos F_{4}[(\bar{\gamma_{3}}^{\sigma}(\gamma_{2};\tilde{\boldsymbol{\alpha}});\hat{\boldsymbol{\alpha}}]:=\bar{\gamma}_{1}^{\hat{\sigma}}(\gamma_{2},\sigma),$ (42)
where $\hat{\boldsymbol{\alpha}}=(\tilde{\alpha}_{2},\tilde{\alpha}_{3},\tilde{\alpha}_{4},\tilde{\alpha}_{1})$. Substituting $\gamma_{3}$ as a function of $\gamma_{2}$, we can still use $\gamma_{2}$ as the reference folding angle in (42). To determine the folding branch $\hat{\sigma}$, we perform the following continuity check for $\tilde{{\bf t}}_{4}$ and $\tilde{{\bf t}}_{55}$:
$\hat{\sigma}(\theta)=\begin{cases}+1,~{}\text{if}~{}{\bf R}_{1}(-\arccos F_{4}(\bar{\gamma_{3}}^{\sigma}(\theta;\tilde{\boldsymbol{\alpha}});\hat{\boldsymbol{\alpha}}))\tilde{{\bf t}}_{55}={\bf R}_{2}(\theta){\bf R}_{3}(\bar{\gamma}_{3}^{\sigma}(\theta))\tilde{{\bf t}}_{4}\\ -1,~{}\text{if}~{}{\bf R}_{1}(\arccos F_{4}(\bar{\gamma_{3}}^{\sigma}(\theta;\tilde{\boldsymbol{\alpha}});\hat{\boldsymbol{\alpha}}))\tilde{{\bf t}}_{55}={\bf R}_{2}(\theta){\bf R}_{3}(\bar{\gamma}_{3}^{\sigma}(\theta))\tilde{{\bf t}}_{4}\end{cases}.$ (43)

### v2

@@ 4 Three-fold intersection
In this section, we study the kinematics of three-fold intersections with patterned piecewise constant director fields satisfying the metric compatibility condition at each interface. Our approach is shown in Fig. 5(a). The reference pattern is metric-compatible at each interface. To compute the actuated configuration, the first step is to cut the reference pattern along the interface ${\bf t}_{1}$ and then deform it by the spontaneous deformation gradient ${\bf U}_{{\bf n}_{i}}$ at each domain. This will form an intermediate configuration in 2D, with a deficit or a surplus. The second step is to deform the intermediate configuration isometrically and stitch the boundaries together to form a non-Euclidean origami with Gaussian curvature at the tip. We use $\alpha_{i}$ and $\tilde{\alpha}_{i}$ to denote the sector angles before and after actuation, respectively.
@@ 4.1 Three-fold origami with angular deficit and surplus
[…]
@@
$\gamma_{2}$ $=$ $\delta\arccos\left(\frac{-\cos\tilde{\alpha}_{3}+\cos\tilde{\alpha}_{1}\cos\tilde{\alpha}_{2}}{\sin\tilde{\alpha}_{1}\sin\tilde{\alpha}_{2}}\right),$
$\gamma_{3}$ $=$ $\delta\operatorname{sign}\left(\frac{\sin\tilde{\alpha}_{1}}{\sin\tilde{\alpha}_{3}}\right)\arccos\left(\frac{-\cos\tilde{\alpha}_{1}+\cos\tilde{\alpha}_{2}\cos\tilde{\alpha}_{3}}{\sin\tilde{\alpha}_{2}\sin\tilde{\alpha}_{3}}\right),$ (13)
for $\tilde{\alpha}_{i}$ in the admissible domain shown in Fig. 6(a), where $\delta\in{\pm 1}$ denotes the two branches.
@@
${\bf R}_{2}(\gamma_{2}){\bf R}_{3}(\gamma_{3})\tilde{{\bf t}}_{4}\cdot\tilde{{\bf t}}_{2}=\tilde{{\bf t}}_{1}\cdot\tilde{{\bf t}}_{2}\Leftrightarrow\cos\gamma_{3}=\frac{-\cos\tilde{\alpha}_{1}+\cos\tilde{\alpha}_{2}\cos\tilde{\alpha}_{3}}{\sin\tilde{\alpha}_{2}\sin\tilde{\alpha}_{3}}.$ (14)
@@
$\gamma_{2}$ $=$ $\delta\arccos\left(\frac{-\cos\tilde{\alpha}_{3}+\cos\tilde{\alpha}_{1}\cos\tilde{\alpha}_{2}}{\sin\tilde{\alpha}_{1}\sin\tilde{\alpha}_{2}}\right),$
$\gamma_{3}$ $=$ $\delta\operatorname{sign}\left(\frac{\sin\tilde{\alpha}_{1}}{\sin\tilde{\alpha}_{3}}\right)\arccos\left(\frac{-\cos\tilde{\alpha}_{1}+\cos\tilde{\alpha}_{2}\cos\tilde{\alpha}_{3}}{\sin\tilde{\alpha}_{2}\sin\tilde{\alpha}_{3}}\right),$ (17)
where $\delta\in\{\pm 1\}$ denotes the two branches. Since $\cos\gamma_{2},\cos\gamma_{3}\in[-1,1]$, we have the admissible domain (Fig. 6(a)) for $\tilde{\alpha}_{1,2,3}$ such that the absolute values of Eqs. (14) and (15) are less than or equal to 1. The symmetric shape of the domain arises from the symmetry of the three sector angles. ∎
Figure 6: (a) The admissible domain of $\tilde{\alpha}_{1}$, $\tilde{\alpha}_{2}$ and $\tilde{\alpha}_{3}$, which ensures the folding angle functions (14) and (15) to have solutions. (b) Notation for a three-fold director field.
@@ 4.2 Connection to director fields
where $\sigma_{i},\tilde{\sigma}_{i}\in\{\pm 1\}$. Equation (18) implies that there are four solutions of ${\bf t}_{2}$ for the prescribed ${\bf n}_{1}$ and ${\bf n}_{2}$. However, these tangents cannot be arbitrarily chosen from the four solutions when patterning a three-fold pattern-they need to obey the following topological condition
@@
An interesting observation from Theorem 4.2 is that, fixing $\sigma_{i}$ and $\tilde{\sigma}_{i}$, the folding angle $\gamma_{i}$ depends only on the directors on its adjacent panels, i.e., ${\bf n}_{i}$ and ${\bf n}_{i-1}$ with the directional angles $\theta_{i}$ and $\theta_{i-1}$. This fact provides more freedom to design connected three-fold intersections. For example, in Fig. 7, the pattern consists of two three-fold intersections. We may keep ${\bf n}_{2}$ and ${\bf n}_{3}$ unchanged and vary the direction of ${\bf n}_{4}$. As a result of Theorem 4.2, the folding angle along the shared crease in the actuated configuration is independent of ${\bf n}_{4}$. The folding angle along the shared crease will remain unchanged. Thus, the actuated configuration containing the two three-fold nodes remains compatible when changing ${\bf n}_{4}$. We will use the basic pattern in Fig. 7 to design more complex patterns in Section 5.
@@ 5 Four-fold intersection
In this section, we study the kinematics of a general non-Euclidean four-fold origami generated by piecewise constant director fields. The kinematics of the four-fold origami (or more general n-fold origami) has been extensively studied (Euclidean or non-Euclidean) by Huffman huffman1976curvature , Santangelo evans2015lattice , van Hecke waitukaitis2015origami ; waitukaitis2020non , Tachi tachi2009generalization , Hull hull1994mathematics , Lang and Howell lang2018rigidly , Guest he2020rigid , Izmestiev izmestiev2017classification , and more, using various methods (e.g. spherical geometry, structural kinematics, mechanisms, etc). However, the continuum-mechanics framework developed in our previous work feng2020designs —which addressed the Euclidean, flat-foldable case—remains worth exploring in the non-Euclidean regime. In particular, we will derive explicit expressions for the folding angle functions and corresponding deformations within the Lagrangian framework, expressed in terms of a reference folding angle defined over the admissible domain. This formulation provides a homotopy that accurately characterizes the evolution of shapes, which is convenient to use in an algorithm.
Again, we deform the intermediate state isometrically to achieve a continuous folded state. The method we use to compute the folding angle functions is partially adapted from our previous work zou2024kinematics . Here, we use a new symmetry approach to make the folding angle functions rigorous and more convenient to use for computation. We also provide a more physically relevant understanding of the “branches” in Fig. 9.
@@ 5.1 Reference domain
Figure 8: (a) Two equivalent intermediate configurations. (b) Two folding branches denoted by $\delta=+1$ and $\delta=-1$.
Let ${\bf R}_{i}(\theta)\in\text{SO(3)}$ be a rotation with axis $\tilde{{\bf t}}_{i}$. Similarly to Eq. (12), the isometric deformation from the intermediate state (Fig. 8(a) top) to the folded state (Fig. 8(b)) is described by
@@ 5.2 Compatibility condition and folding angle functions
The folding angles are restricted in the interval $\pi\leq\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4}\leq\pi$. In the rigidly and flat-foldable case, the four folding angles reach $\pm\pi$ simultaneously, with a reference free folding angle ranging from $\pi$ to $\pi$. In the non-Euclidean case, some folding angles will reach $\pm\pi$ first, making the domain of the reference folding angle not a full interval of $\pi,\pi]$.
@@ 5.2.1 Calculations of $\gamma_{2}$, $\gamma_{3}$ and $\gamma_{4}$
[…]
Here $\delta\in\{\pm\}$ denotes the two branches of solution as shown in Fig. 8(b). Since the function $\bar{\gamma}_{4}^{\delta}(\theta)$ is even and monotonic in $\pi,0]$ and $0,\pi]$, we may characterize the admissible domain of $\theta$ by ${\cal I}^{-}\cup{\cal I}^{+}:=[-\theta_{\max},-\theta_{\min}]\cup[\theta_{\min},\theta_{\max}]$ in the following cases:
@@
[…]
@@
${\bf R}_{3}(\gamma_{3})(({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{4}(\bar{\gamma}_{4}^{\delta}(\gamma_{2}))\tilde{{\bf t}}_{5})=({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}.$ (41)
@@
which is trivially true. Thus, to solve the projected equation (41), we consider the non-degenerate case under $|({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}|\neq 0$. For (41) to hold, $\gamma_{3}=\bar{\gamma}_{3}^{\delta}(\gamma_{2})$ must satisfy
$\cos(\bar{\gamma}_{3}^{\delta}(\gamma_{2}))$ $=$ $\frac{({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{4}(\bar{\gamma}_{4}^{\delta}(\gamma_{2}))\tilde{{\bf t}}_{5}\cdot({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}}{|({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{4}(\bar{\gamma}_{4}^{\delta}(\gamma_{2}))\tilde{{\bf t}}_{5}||({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}|}$ (43)
[…]
$\sin(\bar{\gamma}_{3}^{\delta}(\gamma_{2}))$ $=$ $\text{sign}\left[\tilde{{\bf t}}_{3}\cdot({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{4}(\bar{\gamma}_{4}^{\delta}(\gamma_{2}))\tilde{{\bf t}}_{5}\times({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}\right]\sqrt{1-\cos^{2}(\bar{\gamma}_{3}^{\delta}(\gamma_{2}))}$ (44)
[…]
@@
[…]
@@ 5.2.2 Calculation of $\gamma_{1}$
To solve $\gamma_{1}$, we use the second intermediate domain in Fig. 8(a) such that there exists $\gamma_{1}=\bar{\gamma}_{1}^{\delta}(\theta)$ making the folded configuration continuous across $\tilde{{\bf t}}_{4}$ and $\tilde{{\bf t}}_{55}$. A convenient way to obtain $\gamma_{1}$ is to use the cyclic symmetry of $\tilde{\alpha}_{1},\tilde{\alpha}_{2},\tilde{\alpha}_{3}$ and $\tilde{\alpha}_{4}$. Notice that $\gamma_{4}=\bar{\gamma}_{4}^{\delta}(\gamma_{2})$ shows the ability to solve the folding angle ($\gamma_{4}$) in terms of its opposite folding angle ($\gamma_{2}$). Thus we may solve $\gamma_{1}$ in terms of its opposite folding angle $\gamma_{3}$, by simply cycling $\tilde{\alpha}_{1},\tilde{\alpha}_{2},\tilde{\alpha}_{3},\tilde{\alpha}_{4}$ and then using the the same function form $F_{4}$ in (36) as follows:
$\gamma_{1}=\hat{\delta}\arccos F_{4}[(\bar{\gamma_{3}}^{\delta}(\gamma_{2};\tilde{\boldsymbol{\alpha}});\hat{\boldsymbol{\alpha}}]:=\bar{\gamma}_{1}^{\hat{\delta}}(\gamma_{2},\delta),$ (46)
where $\hat{\boldsymbol{\alpha}}=(\tilde{\alpha}_{2},\tilde{\alpha}_{3},\tilde{\alpha}_{4},\tilde{\alpha}_{1})$. Substituting $\gamma_{3}$ as a function of $\gamma_{2}$, we can still use $\gamma_{2}$ as the reference folding angle in (46). To determine the folding branch $\hat{\delta}$, we perform the following continuity check for $\tilde{{\bf t}}_{4}$ and $\tilde{{\bf t}}_{55}$:
$\hat{\delta}(\theta)=\begin{cases}+1,~\text{if}~{\bf R}_{1}(-\arccos F_{4}(\bar{\gamma_{3}}^{\delta}(\theta;\tilde{\boldsymbol{\alpha}});\hat{\boldsymbol{\alpha}}))\tilde{{\bf t}}_{55}={\bf R}_{2}(\theta){\bf R}_{3}(\bar{\gamma}_{3}^{\delta}(\theta))\tilde{{\bf t}}_{4}\\ -1,~\text{if}~{\bf R}_{1}(\arccos F_{4}(\bar{\gamma_{3}}^{\delta}(\theta;\tilde{\boldsymbol{\alpha}});\hat{\boldsymbol{\alpha}}))\tilde{{\bf t}}_{55}={\bf R}_{2}(\theta){\bf R}_{3}(\bar{\gamma}_{3}^{\delta}(\theta))\tilde{{\bf t}}_{4}\end{cases}.$ (47)